OpenStudy (anonymous):

Need help solving this problem!!!A company will need $75,000 in 6 years for a new addition. To meet this goal, the company deposits money in an account today that pays 6% annual interest compounded quarterly. Find the amount that should be invested to total $75,000 in 6 years. The company should invest $...... Here what I got but don't know if it's right. I=Pe^rt 75,000=P^0.06*6 75,000=0.36

OpenStudy (anonymous):

quarterly compounding not the same as continuous compounding

OpenStudy (anonymous):

\[75,000=P(1+\frac{.06}{4})^{4t}\] solve for t

OpenStudy (anonymous):

sorry that was wrong \[75,000=P(1+\frac{.06}{4})^{24}\] solve for P

OpenStudy (anonymous):

basically in one step \[P=\frac{75,000}{(1.015)^{24}}\]

OpenStudy (anonymous):

Could you show me step by step to this problem, so I know i'm doing it right. I have six more of these two do.

OpenStudy (anonymous):

you are compounding quarterly, i.e. 4 times a year formula for compouning n times a year with interest rate of "r" (as a decimal) is \[P(1+\frac{r}{n})^{nt}\] where P is the principle and t is the time in years

OpenStudy (anonymous):

in your case n = 4, r = .06 and t = 6 so we get a formula of \[P(1+\frac{.06}{4})^{4\times 6}\] or \[P(1.015)^{24}\]

OpenStudy (anonymous):

since you knew you wanted to end with $75,000 and you want P you write \[75,000=P(1.015)^{24}\] and solve for P